The demand function is Q=280000-400p

Where Q equals the number of units demanded and p equals the price in dollars of the total cost of producing Q units of the product is estimated by the function

C=350000+300q+0.0015q2

I) how many units should Q produce in order to maximize its annual profit

Ii) what price should be charged

III) determine the annual profit expected


1
Expert's answer
2022-06-15T05:53:19-0400

Demand Function, "Q = 280\,000-400p"

Inverse Demand Function, "P = \frac{280,000-Q}{400} = 700 - \frac{Q}{400}"


"\text{Total Revenue} = \text{Price} \cdot \text{Quantity} = 700-\frac{Q}{400}\cdot Q = 700Q-\frac{Q^2}{400}"


"\text{Marginal revenue function} = \frac{d}{dQ}(700Q-\frac{Q^2}{400})=700-0.005Q"


"\text{Total Cost function} = 350\,000+300Q+0.0015Q^2"


"\text{Marginal Cost} = \frac{d}{dQ}(350\,000+300Q+0.0015Q^2)=300+0.003Q"


Profit is maximized when Marginal Cost = Marginal Revenue (Price):

"300+0.003Q = 700-0.005Q"

"0.003Q+0.005Q = 700-300"

"0.008Q=400"

"Q=\frac{400}{0.008} = 50\,000"


i) The firm should produce "50\,000" units to maximize its profit


ii) Price that should be charged = Marginal Revenue at "50\,000" units of output = "\$(700 -0.005(50\,000)) = \$(700 - 250) = \$ 450"


iii)At maximum profit condition, "Q = 50\,000"

"\text{Total Revenue} = 700Q-0.005Q^2 = 700 x 50\,000 - 0.005 \cdot 50\,000^2 = \$22\,500\,000"

"\text{Total Cost} = 350\,000+300Q+0.0015Q^2 =" "= 350\,000+300 \cdot 50\,000 + 0.0015 \cdot 50\,000^2 =\$19\,100\,000"

"\text{Annual Profit} = \text{Total Revenue} - \text{Total Cost} = \$22\,500\,000-\$19\,100\,000 = \$3\,400\,000"

Therefore, expected annual profit = "\$3\,400\,000"